# One Model For All: Partial Diffusion for Unified Try-On and Try-Off in Any Pose

### v1

@@ Abstract
Recent diffusion-based approaches have made significant advances in image-based virtual try-on, enabling more realistic and end-to-end garment synthesis. However, most existing methods remain constrained by their reliance on exhibition garments and segmentation masks, as well as their limited ability to handle flexible pose variations. These limitations reduce their practicality in real-world scenarios—for instance, users cannot easily transfer garments worn by one person onto another, and the generated try-on results are typically restricted to the same pose as the reference image. In this paper, we introduce OMFA (One Model For All), a unified diffusion framework for both virtual try-on and try-off that operates without the need for exhibition garments and supports arbitrary poses. For example, OMFA enables removing garments from a source person (try-off) and transferring them onto a target person (try-on), while also allowing the generated target to appear in novel poses—even without access to multi-pose images of that person. OMFA is built upon a novel partial diffusion strategy that selectively applies noise and denoising to individual components of the joint input—such as the garment, the person image, or the face—enabling dynamic subtask control and efficient bidirectional garment-person transformation. The framework is entirely mask-free and requires only a single portrait and a target pose as input, making it well-suited for real-world applications. Additionally, by leveraging SMPL-X–based pose conditioning, OMFA supports multi-view and arbitrary-pose try-on from just one image. Extensive experiments demonstrate that OMFA achieves state-of-the-art results on both try-on and try-off tasks, providing a practical and generalizable solution for virtual garment synthesis.
@@ 1 Introduction
Virtual Try-On (VTON) is the task of generating photorealistic images of a person wearing a desired garment. While recent diffusion-based approaches (Rombach et al. 2022; Zhang, Rao, and Agrawala 2023; Ye et al. 2023; Chen et al. 2024) have significantly advanced this field by improving image realism and enabling end-to-end generation, their practical applicability remains limited. In particular, most existing methods (Choi et al. 2024; Xu et al. 2024; Lee et al. 2022; Kim et al. 2024; Gou et al. 2023) rely on clean garment templates (also referred to as exhibition garments) and segmentation masks to isolate clothing regions. These requirements are often impractical in real-world scenarios—such as mobile shopping platforms or social media—where users typically provide casual images without clean garment inputs or multiple views. Moreover, current methods are generally restricted to preserving the pose of the input image, limiting users’ ability to visualize try-on results in arbitrary or user-specified poses.
@@
To alleviate the impractical need for exhibition garments, some methods (He et al. 2024; Cui et al. 2024) attempt to extract garments directly from images using segmentation algorithms. However, these techniques often suffer from boundary artifacts, occlusion issues, and garment shape distortion, which degrade the final image quality. Alternatively, several recent works (Velioglu et al. 2024; Xarchakos and Koukopoulos 2024) reformulate try-on as a try-off task, aiming to generate a canonical garment representation from a dressed person image. Yet, these approaches treat try-on and try-off as independent tasks, overlooking their inherent duality and failing to provide a unified solution. To address the limitation of pose rigidity, 3D virtual try-on methods (He et al. 2025; Xie et al. 2024) have been explored using parametric body models. However, due to the limited availability of high-quality 3D data compared to 2D imagery, these methods still struggle to produce realistic and high-resolution results.
To overcome these limitations, we propose OMFA (One Model For All), a unified diffusion-based framework for both virtual try-on and try-off. Unlike previous methods, OMFA models garment-person transformations bidirectionally within a single architecture, eliminating the need for garment templates or segmentation masks. At the core of OMFA is a novel partial diffusion mechanism that selectively applies noise and denoising to individual components of the joint input—such as the garment, the person image, or the face—rather than the entire input. This component-wise control enables dynamic subtask modeling and efficient, targeted generation. Our framework is entirely mask-free, inspired by Any2AnyTryOn, and requires only a single portrait and a target pose—making it highly suitable for real-world applications. Furthermore, by incorporating SMPL-X–based (Pavlakos et al. 2019) pose conditioning, OMFA enables multi-view and arbitrary-pose try-on from a single image, supporting immersive and controllable garment synthesis.
@@
  • We introduce OMFA, the first unified framework that jointly performs both virtual try-on and try-off within a single architecture, enabling garment transfer across individuals and directions.
  • We propose a novel partial diffusion mechanism that applies noise selectively to garment-related regions, enabling efficient, localized modeling without reliance on segmentation masks or template garments.
@@ 2 Related Work
Traditional image-based virtual try-on methods primarily rely on Generative Adversarial Networks (GANs), which typically divide the generation process into two separate stages: garment warping and image blending. Deformation techniques such as Thin-Plate Spline (TPS) transformation (Yang et al. 2020; Wang et al. 2018) and optical flow prediction (Lee et al. 2022; Xie et al. 2023) are commonly used to align garments with target body poses, with some methods further employing dense flow fields for finer correspondence. To mitigate misalignment between garments and body structures, approaches such as human parsing normalization (Choi et al. 2021) and knowledge distillation (Ge et al. 2021; Issenhuth, Mary, and Calauzènes 2020) have also been proposed. However, these two-stage pipelines remain sensitive to pose variation and garment complexity, often resulting in artifacts at the fusion boundaries. Moreover, GAN-based models suffer from training instability and limited generalization across diverse identities, garments, and poses.
With the success of diffusion models in image synthesis, a new line of research explores their application in virtual try-on tasks. Early diffusion-based models, such as LaDI-VTON (Morelli et al. 2023) and DCI-VTON (Gou et al. 2023), still depend on explicit garment warping modules. More recent efforts, including TryOnDiffusion (Zhu et al. 2023), move toward one-stage generation by proposing unified architectures (e.g., Parallel-UNet) that learn implicit garment alignment and blending. Several follow-up methods (Kim et al. 2023; Xu et al. 2024; Choi et al. 2024) further refine conditioning mechanisms to better align noisy inputs with denoised outputs. CatVTON (Chong et al. 2024) focuses on computational efficiency by eliminating heavy image encoders, while MV-VTON (Wang et al. 2025) introduces a multi-view setting using both front and back garment views to improve synthesis under non-frontal poses. FitDiT (Jiang et al. 2024) adopts a Diffusion Transformer architecture that incorporates frequency-domain learning to enhance garment texture and fit.
Despite these advances, most diffusion-based try-on models still depend on either clean garment templates or segmentation masks, which are impractical in real-world settings. Furthermore, their synthesis results are typically bound to the input pose, limiting the ability to visualize garments in arbitrary or multi-view poses.
@@
To address the dependence on clean garment templates, several recent works (Velioglu et al. 2024; Xarchakos and Koukopoulos 2024) have explored the task of image-based virtual try-off—i.e., recovering a canonical garment image from a person wearing the clothing. These methods aim to bypass the need for separate garment inputs by extracting garments directly from dressed images. However, most try-off approaches are limited to generating static garment representations and do not integrate naturally with try-on synthesis. They often treat try-on and try-off as separate pipelines, missing the opportunity to unify both tasks into a single generative model. Additionally, try-off models rarely address pose variation or the generation of garments on new identities, limiting their adaptability in dynamic or user-controllable settings.
@@
To improve pose flexibility, 3D-based virtual try-on approaches explicitly (Bridson, Fedkiw, and Anderson 2002; Pons-Moll et al. 2017) or implicitly model human geometry (He et al. 2025). These methods facilitate pose control and multi-view synthesis. However, due to the scarcity of high-quality 3D garment datasets, these models often suffer from low realism in garment texture and shape fidelity. Moreover, 3D fitting pipelines are computationally expensive and less suited for real-time or consumer-facing applications.
@@
In summary, current methods either depend on hard-to-obtain garment inputs, lack flexibility in pose, or treat try-on and try-off as separate problems. In contrast, our proposed OMFA addresses all these limitations by unifying try-on and try-off tasks under a single diffusion-based architecture. OMFA eliminates the need for garment templates and segmentation masks, and enables controllable, pose-aware generation via partial diffusion and SMPL-X conditioning, making it well-suited for practical deployment.
Figure 2: Overview of our proposed OMFA (One Model For All) framework. (a) illustrates the pipeline of person-to-person try-on, including two processes of try-off and try-on in one model. (b) shows the model architecture based on the proposed partial diffusion. The model uses the concatenated multiple conditions as inputs, where the noise are add to the person image (try-on stream) or the garment image (try-off stream). (c) presents the multi-pose try-on support of our framework.
@@ 3 Method
3.1 Preliminaries: Latent Diffusion Models
Diffusion models (Ho, Jain, and Abbeel 2020) are a class of generative models that learn to reverse a forward diffusion process, where noise is progressively added to data. They are trained to denoise a variable, typically starting from pure Gaussian noise $x_{T}$, until a clean data sample $x_{0}$ is produced. In order to reduce computational complexity while generating high-resolution images, Latent Diffusion Models (LDM)(Rombach et al. 2022) operate within the compressed latent space $z=\mathcal{E}(x)$ of a pretrained autoencoder. A prominent example is Stable Diffusion, which employs a network $\epsilon_{\theta}$ to predict the noise in the latent space $z_{t}$. This process is conditioned on the time step $t$ and a condition $c$ like a text embedding, which is integrated into the UNet via cross-attention layers. The model is trained by optimizing the following objective:
$\mathcal{L}_{LDM}=\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,1),c,t}[\|\epsilon-\epsilon_{\theta}(z_{t},c,t)\|_{2}^{2}]$ (1)
Classifier-free guidance (Ho and Salimans 2022) is a widely-used sampling technique that enhances the control over image generation in conditional diffusion models, without needing an extra classifier. This approach requires the diffusion model to be trained on both conditional inputs and unconditional inputs. During the inference step, CFG controls the strength of conditions by linearly combining the conditional and unconditional noise predictions. This guided noise estimate $\tilde{\epsilon}_{\theta}(z_{t},c,t)$, is calculated using a guidance scale $\alpha$ as follows:
$\tilde{\epsilon}_{\theta}(z_{t},c,t)=\epsilon_{\theta}(z_{t},t)+\alpha(\epsilon_{\theta}(z_{t},c,t)-\epsilon_{\theta}(z_{t},t))$ (2)
3.2 Partial Diffusion
To enable component-wise control and improve generation efficiency, we define a partial diffusion process that applies noise and denoising only to selected components of the joint input (composed of the person image, garment, and face). Let $z_{0}=\mathcal{E}(x)$ denote the latent representation of the joint input $x$, and let $M\in\{0,1\}^{\text{shape}(z)}$ be a binary mask indicating which component (e.g., garment or person image) should undergo diffusion.
At diffusion timestep $t$, we construct a partially noised latent as:
$\hat{z}_{t}=M\cdot z_{t}+(1-M)\cdot z_{0},$ (3)
where $z_{t}=\sqrt{\bar{\alpha}_{t}}\cdot z_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon$ is the standard noised latent, and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
During training, the model learns to predict noise only for the diffused component. The training objective is defined as:

### v3

# One Model for All: Unified Try-On and Try-Off in Any Pose via LLM-Inspired Bidirectional Tweedie Diffusion

@@ Abstract
Recent diffusion-based approaches have made significant advances in image-based virtual try-on, enabling more realistic and end-to-end garment synthesis. However, most existing methods remain constrained by their reliance on exhibition garments and segmentation masks, as well as their limited ability to handle flexible pose variations. These limitations reduce their practicality in real-world scenarios—for instance, users cannot easily transfer garments worn by one person onto another, and the generated try-on results are typically restricted to the same pose as the reference image. In this paper, we introduce OMFA (One Model For All), a unified diffusion framework for both virtual try-on and try-off that operates without the need for exhibition garments and supports arbitrary poses. OMFA is inspired by the mask-based paradigm of discrete diffusion language models and unifies try-on and try-off within a bidirectional framework. It is built upon a Bidirectional Tweedie Diffusion process for target-selective denoising in latent space. Instead of imposing lower body constraints, OMFA is an entirely mask-free framework that requires only a single portrait and a target garment as inputs, and is designed to support flexible outfit combinations and cross-person garment transfer, making it better aligned with practical usage scenarios. Additionally, by leveraging SMPL-X–based pose conditioning, OMFA supports multi-view and arbitrary-pose try-on from just one image. Extensive experiments demonstrate that OMFA achieves state-of-the-art results on both try-on and try-off tasks, providing a practical and generalizable solution for virtual garment synthesis. Project page: https://onemodelforall.github.io
@@
11footnotetext: Corresponding author: Guangrun Wang.22footnotetext: These two authors contributed equally and share first authorship.
@@ 1 Introduction
Virtual Try-On (VTON) is the task of generating photorealistic images of a person wearing a desired garment. While recent diffusion-based approaches [33, 50, 48, 3] have significantly advanced this field by improving image realism and enabling end-to-end generation, their applicability remains limited. In particular, most existing methods [5, 46, 22, 21, 11] rely on clean garment templates (also referred to as exhibition garments) and segmentation masks to isolate clothing regions. These requirements are often impractical in real-world scenarios, such as mobile shopping platforms or social media, where users typically provide casual images without clean garment inputs or multiple views. Moreover, current methods are generally restricted to the pose of the input image, limiting users’ ability to visualize try-on results in arbitrary or user-specified poses.
@@
To alleviate the impractical need for exhibition garments, some methods [13, 7] attempt to extract garments directly from images using segmentation algorithms. However, these techniques often suffer from boundary artifacts, occlusion issues, and garment shape distortion, which degrade the final image quality. Alternatively, several recent works [36, 42] reformulate try-on as a try-off task, aiming to generate a canonical garment representation from a dressed person image. Yet, these approaches treat try-on and try-off as independent tasks, overlooking their inherent duality and failing to provide a unified solution. To address the limitation of pose rigidity, 3D virtual try-on methods [14, 43] have been explored using parametric body models. However, due to the limited availability of high-quality 3D data compared to 2D imagery, these methods still struggle to produce realistic and high-resolution results.
To overcome these limitations, we propose OMFA (One Model For All), a unified diffusion-based framework for both virtual try-on and try-off. Unlike previous methods, OMFA models garment-person transformations bidirectionally within a single architecture, eliminating the need for garment templates or segmentation masks. Inspired by the mask-based paradigm of discrete diffusion language models [34, 28], OMFA represents each modality as latent tokens and unifies try-on and try-off as a latent completion problem. At its core, a Bidirectional Tweedie Diffusion process realizes this formulation in continuous latent space by diffusing only the selected target latent while retaining the remaining latents as observed conditions.A shared denoising network is used to handle both generation directions under the same framework, so that the two tasks differ only in the choice of the target latent. Furthermore, by incorporating SMPL-X–based [30] pose conditioning, OMFA enables multi-view and arbitrary-pose try-on from a single image while preserving identity consistency and garment fidelity.
OMFA operates in a practical mask-free setting, requiring only a single portrait image and a target garment image at inference time. Here, “mask-free” means that no segmentation or parsing masks are required during inference. This design is motivated by two practical considerations: the target garment may not always be compatible with the original outfit, and users often prefer flexible combinations of upper and lower garments. Therefore, OMFA does not enforce strict preservation of the non-edited garment or the original pose, allowing more flexible and realistic synthesis.
@@
  • We introduce OMFA, a unified framework that jointly performs both virtual try-on and try-off within a single architecture, enabling bidirectional garment editing without reliance on segmentation masks or template garments.
  • We formulate try-on and try-off as target-selective latent completion under a shared conditional diffusion framework, and instantiate this unified formulation with Bidirectional Tweedie Diffusion.
@@
  • We conduct comprehensive experiments on VITON-HD and Deepfashion-Multimodal datasets, achieving state-of-the-art performance across try-on and try-off in both qualitative and quantitative evaluations.
@@ 2 Related Work
Traditional image-based virtual try-on methods mainly rely on GANs with a two-stage pipeline of garment warping and blending. Although techniques such as TPS [47, 38], optical flow [22, 44], and human parsing priors [4, 9, 18] improve alignment, these methods remain sensitive to pose variation and garment complexity, often producing fusion artifacts and suffering from limited generalization.
With the success of diffusion models in image synthesis, recent virtual try-on research has increasingly shifted toward diffusion-based frameworks. Early methods such as LaDI-VTON [27] and DCI-VTON [11] still rely on explicit garment warping, while later approaches, including TryOnDiffusion [52] and subsequent works [21, 46, 5], move toward one-stage generation with improved conditioning mechanisms for implicit alignment and blending. Other methods further explore efficiency [6], multi-view synthesis [39], and Diffusion Transformer architectures [19]. Despite these advances, most diffusion-based try-on models still depend on clean garment templates or segmentation masks, which are impractical in real-world scenarios, and their results remain constrained by the input pose, limiting arbitrary-pose or multi-view visualization.
@@
To address the dependence on clean garment templates, several recent works [36, 42] have explored the task of image-based virtual try-off—i.e., recovering a canonical garment image from a person wearing the clothing. These methods aim to bypass the need for separate garment inputs by extracting garments directly from dressed images. However, most try-off approaches are limited to generating static garment representations and do not integrate naturally with try-on synthesis. They often treat try-on and try-off as separate pipelines, missing the opportunity to unify both tasks into a single generative model. Additionally, try-off models rarely address pose variation or the generation of garments on new identities, limiting their adaptability in dynamic or user-controllable settings.
@@
To improve pose flexibility, 3D-based virtual try-on approaches explicitly [2, 31] or implicitly model human geometry [14]. These methods facilitate pose control and multi-view synthesis. However, due to the scarcity of high-quality 3D garment datasets, these models often suffer from low realism in garment texture and shape fidelity, partly due to the sim-to-real gap. Moreover, 3D fitting pipelines are computationally expensive and less suited for real-time or consumer-facing applications.
@@
In summary, current methods either depend on hard-to-obtain garment inputs, lack flexibility in pose, or treat try-on and try-off as separate problems. In contrast, our proposed OMFA addresses all these limitations by unifying try-on and try-off tasks under a single diffusion-based architecture. OMFA eliminates the need for garment templates and segmentation masks, and enables controllable, pose-aware generation via a bidirectional Tweedie diffusion mechanism and SMPL-X conditioning, making it well-suited for practical deployment.
Figure 2: Overview of our proposed OMFA (One Model For All) framework. (a) illustrates the pipeline of person-to-person try-on, including two processes of try-off and try-on in one model. (b) depicts a model design based on the Tweedie bidirectional diffusion. The model’s inputs are the latent token sequence, with noise added to the dressed-person image (try-on stream) or the garment image (try-off stream), while the remaining latents are retained as observed conditions. (c) presents the multi-pose try-on support of our framework.
@@ 3 Method
3.1 Overview
We present OMFA (One Model For All), a unified diffusion framework that jointly addresses both virtual try-on and try-off within a single model. Given a person image $I_{p}$, a garment image $I_{g}$, and a dressed person image $I_{d}$, OMFA maps them into a shared latent space, where their semantic dependencies can be modeled under a common conditional generation framework. Within this shared space, try-on and try-off are treated as two task instances of the same latent generation problem: one latent subset is selected as the prediction target, while the remaining subsets are retained as observed conditions.
To instantiate this idea, we formulate generation as conditional latent completion from a masked diffusion perspective, where a selected target latent is recovered from the remaining observed latents (Sec. 3.2). Building on this formulation, we further introduce Bidirectional Tweedie Diffusion in continuous latent space, which provides a unified theoretical interpretation of target-selective denoising for both generation directions (Sec. 3.3). Finally, we incorporate explicit structural guidance through an SMPL-X-based conditioning mechanism, which injects 3D human geometry into the latent generation process and enables pose-controllable and shape-consistent synthesis (Sec. 3.4).
3.2 Conditional Generation as Diffusion Language Modeling
Let $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of a latent diffusion autoencoder (e.g., a VAE). We encode the person image $I_{p}$, garment image $I_{g}$, and dressed image $I_{d}$ into latent variables:
$\mathbf{z}_{p}=\mathcal{E}(I_{p}),\quad\mathbf{z}_{g}=\mathcal{E}(I_{g}),\quad\mathbf{z}_{d}=\mathcal{E}(I_{d}).$ (1)
Rather than treating image domains as fixed inputs and outputs, we formulate these variables as a sequence of generation elements: $\mathcal{Z}=(\mathbf{z}_{p},\mathbf{z}_{g},\mathbf{z}_{d})$. Our method is inspired by diffusion language modeling [28, 34, 45], which excels at generating missing tokens at arbitrary positions within a sequence. Let $\mathcal{Z}^{\mathrm{mask}}$ denote a subset of missing or unobserved latents in the sequence, and $\mathcal{Z}^{\mathrm{obs}}=\mathcal{Z}\setminus\mathcal{Z}^{\mathrm{mask}}$ denote the remaining observed context. Generation is cast as a sequence completion task, where a shared conditional network $f_{\theta}(\cdot)$ learns a unified rule to recover the missing subset:
$p_{\theta}(\mathcal{Z}^{\mathrm{mask}}\mid\mathcal{Z}^{\mathrm{obs}})=f_{\theta}(\mathcal{Z}^{\mathrm{obs}}).$ (2)
This sequence-masking formulation naturally unifies virtual try-on and try-off, as they simply correspond to different missing positions in the sequence. For virtual try-on, the missing element is the dressed-person latent: $\mathcal{Z}^{\mathrm{mask}}=\{\mathbf{z}_{d}\}$, conditioned on $\mathcal{Z}^{\mathrm{obs}}=\{\mathbf{z}_{p},\mathbf{z}_{g}\}$. Conversely, for virtual try-off, the missing element is the garment latent: $\mathcal{Z}^{\mathrm{mask}}=\{\mathbf{z}_{g}\}$, conditioned on $\mathcal{Z}^{\mathrm{obs}}=\{\mathbf{z}_{d}\}$. Therefore, OMFA functions as an arbitrary-position latent completion model. However, unlike traditional discrete diffusion language models which do not follow Tweedie’s formula, we map this sequence-completion concept into a continuous latent space.
3.3 Bidirectional Tweedie Diffusion
Following Sec. 3.2, we formulate OMFA as a Tweedie Diffusion Language Model. Let $m\in\{d,g\}$ denote the selected missing target type in the sequence, where $\mathbf{z}_{0}^{m}$ denotes the clean target latent ($\mathcal{Z}^{\mathrm{mask}}$), and $\mathcal{C}^{m}$ denotes the remaining clean variables used as conditions ($\mathcal{Z}^{\mathrm{obs}}$). We instantiate this masked modeling process by applying continuous Gaussian corruption exclusively to the selected target latent at the missing position, keeping the conditioning variables clean:
$q(\mathbf{z}_{t}^{m}\mid\mathbf{z}_{0}^{m})=\mathcal{N}!\left(\mathbf{z}_{t}^{m};\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}^{m},,(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (3)
where $t$ denotes the diffusion timestep and $\bar{\alpha}_{t}$ denotes the cumulative signal coefficient. Equivalently, the noisy target latent is reparameterized as:
$\mathbf{z}_{t}^{m}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}^{m}+\sqrt{1-\bar{\alpha}_{t}},\epsilon,\qquad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (4)
The fundamental advantage of performing this sequence completion in a continuous domain is that it allows us to mathematically recover the exact posterior mean of the missing content via the Tweedie identity [24, 49]. Under the conditional noisy latent distribution $p_{t}(\mathbf{z}_{t}^{m}\mid\mathcal{C}^{m})$, Tweedie’s formula dictates that the expected value of the clean target is strictly defined by the score of the noisy distribution, $s_{t}^{m}(\mathbf{z}_{t}^{m},\mathcal{C}^{m}):=\nabla_{\mathbf{z}_{t}^{m}}\log p_{t}(\mathbf{z}_{t}^{m}\mid\mathcal{C}^{m})$:
$\mathbb{E}[\mathbf{z}_{0}^{m}\mid\mathbf{z}_{t}^{m},\mathcal{C}^{m}]$ (5)
$=$ $\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{z}_{t}^{m}+(1-\bar{\alpha}t)\nabla{\mathbf{z}_{t}^{m}}\log p_{t}(\mathbf{z}_{t}^{m}\mid\mathcal{C}^{m})\right).$
To compute this score, we rely on Denoising Score Matching [37], which establishes a rigid identity connecting the continuous score field to the expected injected noise:
$\nabla_{\mathbf{z}_{t}^{m}}\log p_{t}(\mathbf{z}_{t}^{m}\mid\mathcal{C}^{m})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\mathbb{E}[\epsilon\mid\mathbf{z}_{t}^{m},\mathcal{C}^{m}].$ (6)
Substituting Eq. (6) directly into Tweedie’s formula (Eq. (5)) yields the exact closed-form identity for recovering the clean target latent entirely in terms of the expected noise:
$\mathbb{E}[\mathbf{z}_{0}^{m}\mid\mathbf{z}_{t}^{m},\mathcal{C}^{m}]=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{z}_{t}^{m}-\sqrt{1-\bar{\alpha}_{t}},\mathbb{E}[\epsilon\mid\mathbf{z}_{t}^{m},\mathcal{C}^{m}]\right).$ (7)
This theoretical formulation directly dictates our practical implementation. To evaluate Eq. (7), we must estimate the conditional expectation of the noise, $\mathbb{E}[\epsilon\mid\mathbf{z}_{t}^{m},\mathcal{C}^{m}]$. By standard statistical decision theory, the optimal parameterized function $\epsilon_{\theta}^{*}$ that exactly matches a conditional expectation is obtained by minimizing the Mean Squared Error (MSE). Therefore, we implement our shared denoising network $\epsilon_{\theta}$ using the standard MSE noise-prediction objective: